# A scanning apparatus to detect the spectral directivity of optically-emissive materials

### v1

@@ 3.3 Small water droplet: off-center pumped
The directivity of the emission spectra for a smaller droplet of wt.% fluorescein/water resting on parafilm was also scanned. The droplet diameter was 2 times smaller than droplet in Section 3.2 which leads to 4 times smaller in size. The pump beam size was approximately same as the large droplet which is m. The pump pulse energy and compact gain region resulted in an above threshold mixture of ASE and fluorescence. Similar to the large droplet described in Section 3.2, the emission intensity increased with the co-latitude as shown in Figure 4(a).
@@
The focused beam was slightly off center in this case which caused an asymmetric emission profile as measured in the azimuthal direction. Figure 4(b) shows emission spectra as a function of the azimuth angle for a fixed co-latitude. The small deviation from the vertical axis centered on the hemispherical droplet had significant effects on the azimuthal dependence of the emission intensity and type of radiative process. The fluorescence and ASE emission intensities were comparable when viewed from an azimuth angle of . As the detector scanned around the sample, the ASE began to dominate the emitted light signal. As with the large droplet, there was a significant amount of uncertainty caused by the shot-to-shot deviations of the OPO pulse energy. Even after considering this uncertainty, there is a clear trend associated with the emission intensity and the azimuthal angle.
4 Discussion
@@
The diffuser-capped LED was initially intended to confirm alignment based on an assumed spherical symmetry for light emission from cosine correction across the hemispherical diffuser cap. In hindsight, the polar angle dependence can be easily explained when assuming a highly directional LED light source directed towards the top of the diffuser cap from the inside. The hot spot identified in the heat map was shown to be caused by the source as opposed to an error in alignment of the apparatus. Additional rotations beyond those shown in Figures 2(a) and 2(b) further confirmed the LED source had an azimuthal dependence on the emission intensity. Thus, the apparatus can be used to characterize electrically-driven light sources.
@@
The high dye concentration in the water droplets was chosen because of the sensitivity of ASE emission. The ASE process occurs when 1) a population inversion density is present and 2) the path length of an emitted beam is long enough to amplify the emitted light. The amplification of the emission when the illumination was nearly cylindrically symmetric resulted in a near azimuthal symmetry for the emission; however, the uncertainty caused by shot-to-shot intensity variations of the OPO pump introduced some noise-like variance between each recorded spectrum which can be observed in Figure 3(a). The directivity was found to be independent of pump polarization.
@@
The characterization for the off-center pumped droplet provided interesting insights. One might expect the ASE and fluorescence intensities to be weighted evenly with respect to each other at all emission angles. Instead we observed different directivities associated with the two emission processes. The results from this last study also illustrate the importance of pump beam characteristics in addition to apparatus alignment when scanning optically pumped light sources.
@@
5 Conclusion
This paper describes an apparatus that scans the spectrum of emitted light as a function of the spherical coordinate angles. Alignment of the apparatus was achieved through the use of a camera which imaged the light source as it swept through a series of angles. The apparatus was found to be well-suited for characterization of LED light sources as well as emissive materials that require optical pumping. Beyond the example materials characterized in Section 3, the directivity of many types of emissive materials can be characterized. Additional examples include random lasers, light sources surrounded by dispersive elements, and complex biological materials stained with fluorescent molecules.
6 Back matter
Funding This material is based upon work supported by the National Science Foundation, Directorate for Mathematical and Physical Sciences (Grant No. 2337595).
Acknowledgment The authors thank Mike Kalman for his role as machinist during the build of the apparatus described in this paper.
Disclosures The presented device is granted with a USA Patent 12203850 on Jan 21 2025.
Data availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.

### v2

@@ IV.3 Small water droplet: off-center pumped
The spectral directive gain density of the emission spectra for a smaller droplet of wt.% fluorescein/water resting on parafilm was also scanned. The droplet in Section IV.2 had twice the diameter of the smaller droplet which equates to four times the cross-sectional area. The pump beam spot was the same size for both droplets and its diameter was measured to be m. The pump pulse energy and compact gain region resulted in an above threshold mixture of ASE and fluorescence. Similar to the large droplet described in Section IV.2, the spectral directive gain density increased with the co-latitude as shown in Figure 4(a).
@@
The focused beam was slightly off center in this case which caused an asymmetric emission profile as measured in the azimuthal direction. Figure 4(b) shows emission spectra as a function of the azimuth angle for a fixed co-latitude. The small deviation from the vertical axis centered on the hemispherical droplet had significant effects on the azimuthal dependence of the emission intensity and type of radiative process. The fluorescence and ASE emission intensities were comparable when viewed from an azimuth angle of . As the detector scanned around the sample, the ASE began to dominate the emitted light signal. As with the large droplet, there was a significant amount of uncertainty caused by the shot-to-shot deviations of the OPO pulse energy. Even after considering this uncertainty, there is a clear trend associated with the spectral directive gain density and the azimuthal angle.
V Discussion
@@
The diffuser-capped LED was initially intended to confirm alignment based on an assumed spherical symmetry for light emission from cosine correction across the hemispherical diffuser cap. In hindsight, the polar angle dependence can be easily explained when assuming a highly directional LED light source directed towards the top of the diffuser cap from the inside. The hot spot identified in the heat map for the scanned-area, spectral directive gain density was shown to be caused by the source as opposed to an error in alignment of the apparatus. Additional rotations beyond those shown in Figures 2(a) and 2(b) further confirmed the LED source had an azimuthal dependence on the emission intensity. Thus, the apparatus can be used to characterize electrically-driven light sources in addition to its initially intended purpose of characterizing gain-doped materials with complex structure functions.
@@
The high dye concentration in the water droplets was chosen because of the sensitivity of ASE emission. The ASE process occurs when 1) a population inversion density is present and 2) the path length of an emitted beam is long enough to amplify the emitted light. The amplification of spontaneous emission when the OPO-pumped droplet was nearly cylindrically symmetric resulted in a near azimuthal symmetry for the emission; however, the uncertainty caused by shot-to-shot intensity variations of the OPO pump introduced some noise-like variance between each recorded spectrum which can be observed in Figure 3(a). The directive gain was found to be independent of pump polarization.
@@
The characterization for the droplet that was pumped off-center provided interesting insights. One might expect the ASE and fluorescence intensities to be weighted evenly with respect to each other at all emission angles. Instead we observed that the directive gains associated with the two emission processes were different from each other as seen in the scanned-area, spectral directive gain density shown in Figure 4. The results from this last study also illustrate the importance of pump beam characteristics in addition to apparatus alignment when scanning optically pumped light sources.
@@
Additional perturbations on the original design are also possible. For example, the reflected and scattered light intensity as a function of wavelength and polar coordinate angles can be measured instead of the emission of a gain-doped sample. In this alternative design, a slewing bearing can be motorized in addition to the plate attached to the borehole motor which allows for a source (such as a collimated flat-spectrum LED) and a detector to rotate in the azimuthal direction independently. The source and detector can then have two separate motors to adjust the polar angle for each one. From such an alternative design, the reflected and scattered light could be measured as a function of source position, detector position, and wavelength, i.e., the detected light intensity would follow as .
VI Conclusion
This article describes an apparatus that scans the spectrum of emitted light as a function of the spherical coordinate angles. Alignment of the apparatus was achieved through the use of a camera which imaged the light source as it swept through a series of angles. The apparatus was found to be well-suited for characterization of LED light sources as well as emissive materials that require optical pumping. Beyond the example materials characterized in Section IV, the scanned-area, spectral directive gain density of many types of emissive materials can be characterized. Additional examples include random lasers, light sources surrounded by dispersive elements, and complex biological materials stained with fluorescent molecules.
Funding: This material is based upon work supported by the National Science Foundation, Directorate for Mathematical and Physical Sciences (Grant No. 2337595).
Acknowledgment: The authors thank Mike Kalman for his role as machinist during the build of the apparatus described in this paper.
Disclosures: The presented device was granted a US Patent 12203850 on Jan 21 2025.
Data availability: Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.